Full conversion of 12 hour and 30 minutes into seconds........

Grade:12th pass

10 Answers

venky
8 years ago
4500 seconds
KARTHIK
8 years ago
4500 SECONDS
pa1
8 years ago
45000 sec
Gaddam Chethan
8 years ago
1hour=60 minutes, 1minutes=60 seconds. 12*60=720 minutes. 720+30=750minutes.750*60=45000 seconds. it means (12:30= 45000 seconds)
Nitesh
6 years ago
First change hour into minute Since 1 hour = 60 minuteSo, 12 hour=12×60 minute=》12 hour = 720 minute Total minute= 720+30=750 minuteNow convert minute into seconds 1 minute= 60 seconds=》750 minute = 750×60 seconds=》750 minute=45000 seconds hence 45000 seconds is the answer
C.Vishaal
6 years ago
Conversion of 12 hours and 30 minutes into seconds can be done by:
 
1 hr = 60 min
1 min = 60 sec.
 
Therefore 60 min = 1hr = 60 x 60 secs = 3,600 secs.
 
1 hr = 3,600 secs.
12hrs = 3,600 x 12 = 43,200 secs...................................  (1)
 
1 min = 60 secs
30 mins = 30 x 60 = 1,800 sec.........................................(2)
 
From (1) and (2), we get that 12hrs + 30 mins = 43,200 + 1,800 = 45,000 seconds .
 
Hope this answer helps.
 
Regards
C.Vishaal
Priyansh
6 years ago
45000 seconds bro.............................................................................................................................................
SR Roy
6 years ago
1 hour = 60 min
12 hours = 60X12 min = 720min
now, 720+30min = 750min
1min= 60sec
750min = 750X60sec= 45000sec.
 
Ans. 45000sec
regards,
Suha
Varsha Shivanand K
5 years ago
1 Hour=60 minutes
12 hours=12×60 minutes=720 minutes
720mins+30 mins=750 minutes
1 min = 60 seconds
750 min=750×60 sec = 45000 seconds.
 
Answered by Varsha Shivanand K
Sayantan Garai
5 years ago
12h30min
=(12×3600)+(30×60) s
=43200+1800 s
=45000 s (Answer).
--------------------------------------------------------------